OpenStudy (anonymous):

A. (100 + 82 + 70 + 72 + 80 + x)/6 ≤ 80 (404 + x)/6 ≤ 80 (404 + x)/6 X 6 ≤ 80 X 6 404 + x ≤ 480 404 + x - 404 ≤ 480 - 404 x = 76 B. The minimum grade that Sam must earn in order to keep a B average in math is 76%. Is this correct? Or does it need to be an equals sign? I need it to be an algebraic inequality though.

OpenStudy (anonymous):

x≤76 Sam needs at least a 76, so he could have more.

OpenStudy (anonymous):

Why is it less than or equal to? Doesn't he need more? Since he needs a B average?

OpenStudy (anonymous):

You're right. It does need to be greater than or equal to. Your work, however, shows that hs average needs to be less than or equal to 80. I apologize.

OpenStudy (anonymous):

ohhhh can you tell me how to fix it?

myininaya (myininaya):

\[80 \le \frac{100+82+70+80+x}{6} \le 89.99999999\]

OpenStudy (anonymous):

Flip the sign in your first line of work, and in all subsequent lines.

myininaya (myininaya):

\[80 \le \frac{100+82+70+80+x}{6} < 90\] this is one is better than mine other inequality

OpenStudy (anonymous):

so, 80 should be like this:\[80\ge\] instead?

OpenStudy (anonymous):

Like this: 80 ≤ (100 + 82 + 70 + 72 + 80 + x)/6

OpenStudy (anonymous):

ohhhh I see

OpenStudy (anonymous):

should it be like this? x \[\le\] 76?

myininaya (myininaya):

oh i missed one number 72in my inequalities above \[480 \le 404+x <540\] \[76 \le x <136\] to make a B x needs to be between =76 and 136 (but part of this range is impossible) so really to make a B you need between =76 and =100 the minimum you need to make is a 76 to make a B

OpenStudy (anonymous):

Ohhh I see, so is this right? 80 ≤ (100 + 82 + 70 + 72 + 80 + x)/6 < 90 80 ≤ (404 + x)/6 < 90 80 X 6 ≤(404 + x)/6 X 6 < 90 X 6 480 ≤ 404 + x < 540 480 - 404 ≤404 + x - 404 < 540 - 404 76 ≤ x < 136 x = 76

OpenStudy (anonymous):

x ≤ 76 at the end right?

myininaya (myininaya):

no x greater than or equal to 76 not less than of equal to

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

Thank you for your help!